OpenStudy (anonymous):

Solve please! \[\left| 3x(4x-5) \right|=3\]

OpenStudy (anonymous):

well clean that up

OpenStudy (anonymous):

|12x^2-15x|=3

OpenStudy (anonymous):

im in that step...

OpenStudy (anonymous):

12x^2-15=3 or 12x^2-15=-3

OpenStudy (anonymous):

solve for each case

OpenStudy (anonymous):

would it be (2,-2)??

OpenStudy (anonymous):

no

OpenStudy (anonymous):

agh

OpenStudy (anonymous):

(2,8) !! ?

OpenStudy (anonymous):

yea it is xD

OpenStudy (anonymous):

12x^2-15=3 12x^2=18 x^2 = 3/2